OpenStudy (anonymous):

x minus 2 over 5 equals x minus 4 over 2. solve for x how do i do this problem please explain in detail

myininaya (myininaya):

\[\frac{x-2}{5}=\frac{x-4}{2}\]

OpenStudy (saifoo.khan):

cross multiply.

myininaya (myininaya):

cross multiply!

OpenStudy (saifoo.khan):

2x - 4= 5x - 20

myininaya (myininaya):

\[2(x-2)=5(x-4)\] you can solve this :)

OpenStudy (anonymous):

then put all the x's on oneside and the numbers on the other

OpenStudy (saifoo.khan):

3x = 16

OpenStudy (anonymous):

x=5 1/3

OpenStudy (anonymous):

saifoo.khan great job

myininaya (myininaya):

saifoo is right

OpenStudy (anonymous):

got got got it thanks a bunch!